# A model of full thermodynamic stabilization of nanocrystalline alloys

### v2

@@ 1 Introduction
Over the past decades, nanocrystalline (NC) materials have attracted much attention in both research and technological applications [1]. Some mechanical and physical properties of NC materials are superior to those of their conventional, coarse-grained counterparts. However, NC materials are thermodynamically unstable because grain boundaries (GBs) possess an excess free energy relative to the perfect crystalline lattice. This excess free energy causes extensive grain growth at elevated temperatures. Once the grains grow in size well above the nanoscale, the superior properties of the NC material are lost.
@@
In the 1990s, Weissmuller [45, 46] and later others [47, 48, 49, 50, 51, 52, 3, 4, 5, 6, 7, 8, 9, 10, 11] discussed the possibility of decreasing the GB free energy until the total free energy of the polycrystal reaches a minimum at a finite grain size.***More precisely, the free energy $F$ is minimized with respect to the fraction $f$ of GB sites: $dF/df=0$. It is assumed that all GBs have the same width $\delta$ and uniform properties. Assuming also that $f\propto\delta/D$, where $D>0$ is the average grain size, the stability condition becomes $dF/dD=0$. In the minimum-free-energy state, the driving force for grain growth vanishes, and the NC material becomes fully stabilized. Thermodynamic analysis within the uniform boundary model predicts that in order to achieve the fully stabilized state, the GB free energy $\gamma$ must converge to a zero value when the system arrives at the equilibrium state [53].
@@ 2 Model formulation
We consider a simple square lattice forming a rectangular simulation block with periodic boundary conditions. The block contains $N$ sites enumerated by the index $k=1,2,...,N$. Each site is interpreted as a small crystallite capable of adopting one of $q$ distinct lattice orientations labeled by index $\sigma=1,2,...,q$.
@@
Under certain conditions discussed later, the system breaks into domains composed of sites with the same orientation, which we interpret as grains separated by GBs. An example of a polycrystalline structure is shown in Fig. 1(a). Inside the grains, most sites have $n_{k}=0$ (perfect lattice) and occasionally $n_{k}=4$ (isolated inclusion with a “wrong” orientation) or $n_{k}=1$ (neighbors of the isolated inclusion). In contrast, most sites in GB regions have $n_{k}=2\pm 1$. To identify GBs and track their motion, we introduce the GB locator function
@@
We consider two types of transitions, which we call flips and jumps. In a flip transition, a chosen site $k$ switches its current orientation to one of the remaining $q-1$ orientations with $\sigma_{l}\neq\sigma_{k}$. If a group of sites flips into the same orientation on one side of a GB, this can cause GB migration. In a jump event, a chosen solute atom jumps to a nearest-neighbor site if the latter is unoccupied. Such jumps represent solute diffusion. The flip and jump transitions are assigned unbiased barriers $\varepsilon_{0}^{g}$ and $\varepsilon_{0}^{s}$, respectively. The flip barrier $\varepsilon_{0}^{g}$ is the same for all sites and is related to the activation energy of GB migration. The jump barrier $\varepsilon_{0}^{s}$ represents the activation energy of solute diffusion. In the present model, $\varepsilon_{0}^{s}$ is assigned different values in GBs and inside the grains. Specifically, the unbiased barrier of a jump is
@@
where $\varepsilon_{00}^{s}$ is the jump barrier inside the grains and $n_{k,l}$ is the average number of misoriented neighbors in the initial ($k$) and final ($l$) positions of the solute atom. The term in square brackets in Eq.(10) lowers the barrier for jumps within GB regions. Since in GB environments $\phi(n_{k,l})\approx 1$, the barrier becomes approximately $\varepsilon_{0}^{s}\approx\varepsilon_{00}^{s}\left[1-\eta\right]$. The parameter $10\leq\eta<$ 1 controls the decrease in the diffusion barrier leading to accelerated GB diffusion (“short-circuit” diffusion [56, 57]) relative to the perfect lattice. In the simulations, we usually take $\eta=1/2$ to match the experimentally established correlation $\varepsilon_{0}^{s}\approx\varepsilon_{00}^{s}/2$ between the activation energies of GB and lattice diffusion [56]. Note that a solute jump is generally accompanied by a change in the total energy ($E_{ij}\neq 0$). As a result, the actual jump barrier $\varepsilon_{ij}$ is biased relative to $\varepsilon_{0}^{s}$ according to Eq.(9).
@@
Two features distinguish our KMC algorithm from previous Monte Carlo simulations using the Ising and Potts models with solutes [59, 60, 10, 5, 61, 62]. Firstly, we replace the Metropolis scheme with the probability calculations using the actual transition barriers. This scheme is more suitable for the modeling of solute diffusion and GB migration than the Metropolis algorithm. The latter correctly samples canonical fluctuations in equilibrium states but the path toward equilibrium is not guaranteed to be physically meaningful. Secondly, previous simulations within the Potts model only allowed flips to one of the neighboring orientations. In contrast, we allow the current orientation to switch to any of the alternative $q-1$ orientations. This imposes fewer constraints on the system evolution and captures processes such as new grain nucleation and orientational fluctuations in single crystals as well as disordered, liquid-like structures.
@@ 3.2 The alloy phase diagrams
In addition to the structural distinctions, the supercooled liquid displays much slower dynamics compared to the high-temperature liquid. For example, Fig. 5(c) shows the Arrhenius plot of the orientation switching rate $r$ as a function of temperature at a fixed chemical composition. To calculate $r$, we keep track of the orientation changes at all sites during an equilibrium KMC run. At each site, we compute the time interval $\Delta t$ between the moment when the site switched to its current orientation and the moment when it switches to a new orientation. After the switch, we reset the clock and start counting the time to the next switch, and so on. The value $\overline{\Delta t}$ averaged over a long KMC run and all sites represents the time scale on which the sites preserve their orientation. Accordingly, $r=1/\overline{\Delta t}$ characterizes the rate of orientational changes in the liquid, which is a measure of its intrinsic dynamics. The plot of $\log r$ versus $1/T$ shown in Fig. 5(c) displays a non-Arrhenius behavior. The increase in $r$ with temperature accelerates above $T\approx 0.35$, which is approximately where the structural changes in the liquid are observed. We associate the structural and dynamical changes in the liquid at $T\approx 0.35$ with the formation of supercooled liquid. The temperature of this transition depends on the chemical composition. Because this transition is continuous, its position on the phase diagram is approximate and is shown by a dashed line.
@@ 4 Discussion and conclusions
We proposed a simple 2D model of a polycrystalline alloy combining the Potts model for grain orientations with a lattice-gas model for solute thermodynamics and diffusion. The model couples the evolution of the crystal orientations (and thus GB migration) to thermodynamic properties of the alloy and the solute diffusion. The model recognizes that the thermodynamic parameters and the solute diffusion coefficients are generally different inside the grains and at GBs. The model is solved by KMC simulation with nonlinear transition barriers. We have applied this model to investigate the problem of polycrystalline stability, but it has a much broader applicability. It can be used to study other processes, such as solute drag and the impact of solute segregation on GB diffusion.
The proposed model does not represent any particular material and is intended for a generic analysis of thermodynamics and kinetics of polycrystalline alloys. On the other hand, the model goes beyond the uniform boundary approximation, automatically includes nucleation processes due to its stochastic nature, and captures the diffusive time scales not accessible by material-specific methods such as molecular dynamics simulations.
@@
Although this work does not point to a particular alloy system in which a nano-stabilization can be achieved, we believe that the results can be useful as a general guide in this pursuit. In particular, they provide a glimpse into what a fully stabilized NC state would look like if it were achieved experimentally in the future.

### v3

@@ 1 Introduction
Over the past decades, nanocrystalline (NC) materials have attracted much attention in both research and technological applications [1, 2]. Some mechanical and physical properties of NC materials are superior to those of their conventional, coarse-grained counterparts. However, NC materials are thermodynamically unstable because grain boundaries (GBs) possess an excess free energy relative to the perfect crystalline lattice. This excess free energy causes extensive grain growth at elevated temperatures. Once the grains grow in size well above the nanoscale, the superior properties of the NC material are lost.
@@
In the 1990s, Weissmuller [46, 47] and later others [48, 49, 50, 51, 52, 53, 4, 5, 6, 7, 8, 9, 10, 11, 12, 54] discussed the possibility of decreasing the GB free energy until the total free energy of the polycrystal reaches a minimum at a finite grain size.***More precisely, the free energy $F$ is minimized with respect to the fraction $f$ of GB sites: $dF/df=0$. It is assumed that all GBs have the same width $\delta$ and uniform properties. Assuming also that $f\propto\delta/D$, where $D>0$ is the average grain size, the stability condition becomes $dF/dD=0$. In the minimum-free-energy state, the driving force for grain growth vanishes, and the NC material becomes fully stabilized. Thermodynamic analysis within the uniform boundary model predicts that in order to achieve the fully stabilized state, the GB free energy $\gamma$ must converge to a zero value when the system arrives at the equilibrium state [55].
@@ 2 Model formulation
We consider a simple square lattice forming a rectangular simulation block with periodic boundary conditions. The block contains $N$ sites enumerated by the index k=1,2,,Nk=1,2,...,N. Each site is interpreted as a small crystallite capable of adopting one of $q$ distinct lattice orientations labeled by index $\sigma=1,2,...,q$.
@@
Under certain conditions discussed later, the system breaks into domains composed of sites with the same orientation, which we interpret as grains separated by GBs. An example of a polycrystalline structure is shown in Fig. 1(a). In this and other figures in this paper, the colors represent different crystal orientations corresponding to different $\sigma$-numbers. Inside the grains, most sites have $n_{k}=0$ (perfect lattice) and occasionally $n_{k}=4$ (isolated inclusion with a “wrong” orientation) or $n_{k}=1$ (neighbors of the isolated inclusion). In contrast, most sites in GB regions have $n_{k}=2\pm 1$. To identify GBs and track their motion, we introduce the GB locator function
@@
We consider two types of transitions, which we call flips and jumps. In a flip transition, a chosen site $k$ switches its current orientation to one of the remaining q1q-1 orientations with $\sigma_{l}\neq\sigma_{k}$. If a group of sites flips into the same orientation on one side of a GB, this can cause GB migration. In a jump event, a chosen solute atom jumps to a nearest-neighbor site if the latter is unoccupied. Such jumps represent solute diffusion. The flip and jump transitions are assigned unbiased barriers $\varepsilon_{0}^{g}$ and $\varepsilon_{0}^{s}$, respectively. The flip barrier $\varepsilon_{0}^{g}$ is the same for all sites and is related to the activation energy of GB migration. The jump barrier $\varepsilon_{0}^{s}$ represents the activation energy of solute diffusion. In the present model, $\varepsilon_{0}^{s}$ is assigned different values in GBs and inside the grains. Specifically, the unbiased barrier of a jump is
@@
where $\varepsilon_{00}^{s}$ is the jump barrier inside the grains and $n_{k,l}$ is the average number of misoriented neighbors in the initial ($k$) and final ($l$) positions of the solute atom. The term in square brackets in Eq.(10) lowers the barrier for jumps within GB regions. Since in GB environments $\phi(n_{k,l})\approx 1$, the barrier becomes approximately $\varepsilon_{0}^{s}\approx\varepsilon_{00}^{s}\left[1-\eta\right]$. The parameter $0\leq\eta<1$ controls the decrease in the diffusion barrier leading to accelerated GB diffusion (“short-circuit” diffusion [58, 59]) relative to the perfect lattice. In the simulations, we usually take $\eta=1/2$ to match the experimentally established correlation $\varepsilon_{0}^{s}\approx\varepsilon_{00}^{s}/2$ between the activation energies of GB and lattice diffusion [58]. Note that a solute jump is generally accompanied by a change in the total energy ($E_{ij}\neq 0$). As a result, the actual jump barrier $\varepsilon_{ij}$ is biased relative to $\varepsilon_{0}^{s}$ according to Eq.(9).
@@
Two features distinguish our KMC algorithm from previous Monte Carlo simulations using the Ising and Potts models with solutes [61, 62, 11, 6, 63, 64]. Firstly, we replace the Metropolis scheme with the probability calculations using the actual transition barriers. This scheme is more suitable for the modeling of solute diffusion and GB migration than the Metropolis algorithm. The latter correctly samples canonical fluctuations in equilibrium states but the path toward equilibrium is not guaranteed to be physically meaningful. Secondly, previous simulations within the Potts model only allowed flips to one of the neighboring orientations. In contrast, we allow the current orientation to switch to any of the alternative q1q-1 orientations. This imposes fewer constraints on the system evolution and captures processes such as new grain nucleation and orientational fluctuations in single crystals as well as disordered, liquid-like structures.
@@ 3.2 The alloy phase diagrams
In addition to the structural distinctions, the supercooled liquid displays much slower dynamics compared to the high-temperature liquid. For example, Fig. 5(c) shows the Arrhenius plot of the orientation switching rate $r$ as a function of temperature at a fixed chemical composition. To calculate $r$, we keep track of the orientation changes at all sites during an equilibrium KMC run. At each site, we compute the time interval $\Delta t$ between the moment when the site switched to its current orientation and the moment when it switches to a new orientation. After the switch, we reset the clock and start counting the time to the next switch, and so on. The value $\overline{\Delta t}$ averaged over a long KMC run and all sites represents the time scale on which the sites preserve their orientation. Accordingly, $r=1/\overline{\Delta t}$ characterizes the rate of orientational changes in the liquid, which is a measure of its intrinsic dynamics. The plot of $\log r$ versus $1/T$ shown in Fig. 5(c) displays a non-Arrhenius behavior. The increase in $r$ with temperature accelerates above $T\approx 0.35$, which is approximately where the structural changes in the liquid are observed. We associate the structural and dynamical changes in the liquid at $T\approx 0.35$ with the formation of supercooled liquid. The temperature of this transition depends on the chemical composition. Because this transition is continuous, its position on the phase diagram is approximate and is shown by a dashed line. This line separates the stable polycrystalline state from a single crystal with inclusions assuming a threshold grain size of 12 sites.
@@ 4 Discussion and conclusions
We proposed a simple 2D model of a polycrystalline alloy combining the Potts model for grain orientations with a lattice-gas model for solute thermodynamics and diffusion. The model couples the evolution of the crystal orientations (and thus GB migration) to thermodynamic properties of the alloy and the solute diffusion. It recognizes that the thermodynamic parameters and the solute diffusion coefficients are generally different inside the grains and at GBs. The model goes beyond the uniform boundary approximation, automatically includes nucleation processes due to its stochastic nature, and captures the diffusive time scales not accessible by material-specific methods such as molecular dynamics simulations. The model is solved by KMC simulation with nonlinear transition barriers. We have applied this model to investigate the problem of polycrystalline stability, but it has a much broader applicability. It can be used to study other processes, such as solute drag and the impact of solute segregation on GB diffusion.
@@
Although the proposed model does not represent any particular material and is intended for a generic analysis of thermodynamics and kinetics of polycrystalline alloys, it was important to map the domain of the model parameters explored in our simulations onto thermodynamic properties of real alloy systems. In the Supplementary Information file accompanying this article, we have expressed the model parameters through the experimentally accessible properties such as the heat of solution in the dilute limit, $H_{\mathrm{mix}}$, the GB segregation energy $E_{s}$, the GB free energy in the pure solvent material $\gamma$, and the regular solution parameter $\Omega^{\prime}$ inside the GBs (which is generally different from $\Omega$ within the grains). The estimates are based on two reasonable assumptions: the area per atom in the GB core is on the order of 1 nm2, and $\gamma$ is about 1 J/m2.
Table 2 summarizes the results for three representative parameter sets corresponding to the phase diagrams shown in Figures 3, 6, and 7. The predicted magnitudes of $E_{s}$, $\Omega$, $\Omega^{\prime}$, and $H_{\mathrm{mix}}$ are well within the typical ranges reported in previous publications [11, 5, 6, 68, 7, 8, 9]. A notable feature of the present simulations is that the stable polycrystalline structure was found in alloys with a negative heat of mixing, corresponding to the ordering trend (Table 2 and Fig. 7). Most of the previous calculations [11, 5, 6, 68, 7, 8, 9] considered phase-separating alloys with $H_{\mathrm{mix}}>0$. However, recent publications [68, 7] extended the previous analyses to alloys with $H_{\mathrm{mix}}<0$ and still reported stable nanocrystalline states. For example, Figure 6 in [68] presents a diagram showing a nanocrystalline stability domain at $\Omega^{\prime}<\Omega<0$. Furthermore, the values of $\Omega$ and $\Omega^{\prime}$ for the representative point g in this domain are comparable to the respective values in Table 2. A more detailed quantitative comparison with previous publications [11, 5, 6, 68, 7, 8, 9] is complicated due to intrinsic differences between models (see Supplementary Information for more detail), but the qualitative agreement indicates that the model parameters investigated in this paper are practically relevant.
Due to the 2D constraint and the generic character of this model, the results do not allow us to pinpoint a particular alloy system in which a nano-stabilization can be achieved. However, we believe that the results can be useful as a general guide in this pursuit. In particular, they provide a glimpse into what a fully stabilized NC state would look like if it were achieved experimentally in the future.
@@ Acknowledgements
5 Mapping of Potts model onto solution thermodynamics
Consider a binary solid solution of components A and B with nearest-neighbor bond energies $\varepsilon_{AA}$, $\varepsilon_{AB}$, and $\varepsilon_{BB}$. The total energy of the solution can be written in our notation as
$E_{c}$ $=$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}\xi_{k}\left[\xi_{l}\varepsilon_{BB}+(1-\xi_{l})\varepsilon_{AB}\right]$ (1)
$+$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}(1-\xi_{k})\left[\xi_{l}\varepsilon_{AB}+(1-\xi_{l})\varepsilon_{AA}\right]$
$=$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}\varepsilon_{AA}+\sum_{k}\sum_{(k,l)}\xi_{k}(\varepsilon_{AB}-\varepsilon_{AA})$
$+$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}\xi_{k}\xi_{l}(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB}).$
Here, $\xi_{k}=1$ if the site (cell) $k$ is occupied by an atom of species $B$ and $\xi_{k}=0$ otherwise. The symbol $(kl)$ indicates the summation over all nearest-neighbors $l$ of a site $k$.
In addition to the occupation of the solutes, the sites are characterized by orientations. To include the orientational energy, we introduce the wrong bond parameter $m_{kl}$ such that $m_{kl}=1$ if the neighboring sites $k$ and $l$ have different orientations and $m_{kl}=0$ otherwise. Eq.(1) is modified by replacing $\varepsilon_{AA}$, $\varepsilon_{AB}$, and $\varepsilon_{BB}$ with, respectively,
$\varepsilon_{AA}+m_{kl}(\varepsilon_{AA}^{\prime}-\varepsilon_{AA}),\>\>\>\varepsilon_{AB}+m_{kl}(\varepsilon_{AB}^{\prime}-\varepsilon_{AB}),\>\>\>\varepsilon_{BB}+m_{kl}(\varepsilon_{BB}^{\prime}-\varepsilon_{BB}).$
Here, the primed energies represent distorted bonds between misoriented sites. Because Eq.(1) is linear in the bond energies, it can be immediately rewritten in the form
$E$ $=$ $E_{c}+\dfrac{1}{2}\sum_{k}\sum_{(k,l)}m_{kl}(\varepsilon_{AA}^{\prime}-\varepsilon_{AA})+\sum_{k}\sum_{(k,l)}\xi_{k}m_{kl}\left[(\varepsilon_{AB}^{\prime}-\varepsilon_{AA}^{\prime})-(\varepsilon_{AB}-\varepsilon_{AA})\right]$ (2)
$+$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}m_{kl}\xi_{k}\xi_{l}\left[(\varepsilon_{BB}^{\prime}+\varepsilon_{AA}^{\prime}-2\varepsilon_{AB}^{\prime})-(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB}\right].$
This equation can be simplified as follows:
$E$ $=$ $E_{c}+\dfrac{1}{2}\sum_{k}(\varepsilon_{AA}^{\prime}-\varepsilon_{AA})n_{k}+\sum_{k}\left[(\varepsilon_{AB}^{\prime}-\varepsilon_{AA}^{\prime})-(\varepsilon_{AB}-\varepsilon_{AA})\right]\xi_{k}n_{k}$ (3)
$+$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}m_{kl}\xi_{k}\xi_{l}\left[(\varepsilon_{BB}^{\prime}+\varepsilon_{AA}^{\prime}-2\varepsilon_{AB}^{\prime})-(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB}\right],$
where $n_{k}=\sum_{(k,l)}m_{kl}$ is the number of unlike (wrong) nearest-neighbors of the site $k$.
In the Potts model with solutes introduced in the main text of the paper, the total energy is postulated in the form
$E$ $=$ $\sum_{k}J_{gg}n_{k}$ (4)
$+$ $\sum_{k}J_{s}\xi_{k}+\sum_{k}J_{sg}\xi_{k}\phi(n_{k})$
$+$ $\dfrac{1}{2}\sum_{k}\sum_{(kl)}\xi_{k}\xi_{l}J_{ss}+\dfrac{1}{2}\sum_{k}\sum_{(kl)}\xi_{k}\xi_{l}J_{ssg}\phi(n_{k,l}).$
Here, $\phi(n_{k})$ is the grain boundary locator function defined in the main text, $n_{k,l}$ is the average number of wrong bonds of the sites $k$ and $l$, and $J_{s}$ is the energy on site of a solute atom. The remaining $J$ factors in Eq.(4) characterize the interaction energies between misoriented sites ($J_{gg}$), between the solute atoms and grain boundaries ($J_{sg}$), and between solute atoms inside the grains ($J_{ss}$) and at grain boundaries ($J_{ssg}$).
Let us rewrite Eq.(3) in a form closest to Eq.(4):
$E$ $=$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}\varepsilon_{AA}+\dfrac{1}{2}\sum_{k}(\varepsilon_{AA}^{\prime}-\varepsilon_{AA})n_{k}$ (5)
$+$ $\sum_{k}\sum_{(k,l)}\xi_{k}(\varepsilon_{AB}-\varepsilon_{AA})+\sum_{k}\left[(\varepsilon_{AB}^{\prime}-\varepsilon_{AA}^{\prime})-(\varepsilon_{AB}-\varepsilon_{AA})\right]\xi_{k}n_{k}$
$+$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}\xi_{k}\xi_{l}(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB})$
$+$ $\dfrac{1}{2}\sum_{k}\sum_{(k,l)}m_{kl}\xi_{k}\xi_{l}\left[(\varepsilon_{BB}^{\prime}+\varepsilon_{AA}^{\prime}-2\varepsilon_{AB}^{\prime})-(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB})\right].$
Our goal is to establish relationships between the $J$-factors in Eq.(4) and the bond energies $\varepsilon_{ij}$ in Eq.(5) by mapping the two equations to each other term by term.
The first term in Eq.(5) is the energy of the pure solvent A. This term is not present in Eq.(4). Since this term is a constant, we can drop it without loss of generality by assuming that $\varepsilon_{AA}=0$. The second term in Eq.(5) maps onto the solute-free energy:
$\dfrac{1}{2}\sum_{k}(\varepsilon_{AA}^{\prime}-\varepsilon_{AA})n_{k}=\sum_{k}J_{gg}n_{k},$
from which
$J_{gg}=\dfrac{1}{2}(\varepsilon_{AA}^{\prime}-\varepsilon_{AA}).$ (6)
The third term is
$\sum_{k}\sum_{(k,l)}\xi_{k}(\varepsilon_{AB}-\varepsilon_{AA})=\sum_{k}\xi_{k}z(\varepsilon_{AB}-\varepsilon_{AA}),$ (7)
where $z$ is the number of closest neighbors ($z=4$ in this model). Here, $z(\varepsilon_{AB}-\varepsilon_{AA})$ has the meaning of the energy cost of replacing an atom A with an atom B in the pure-A perfect lattice. In this process, we break $z$ AA bonds and create $z$ AB bonds. The term in Eq.(7) maps on the term $\sum_{k}J_{s}\xi_{k}$ appearing Eq.(4). It follows that
$J_{s}=z(\varepsilon_{AB}-\varepsilon_{AA}).$ (8)
The next term,
$\sum_{k}\left[(\varepsilon_{AB}^{\prime}-\varepsilon_{AA}^{\prime})-(\varepsilon_{AB}-\varepsilon_{AA})\right]\xi_{k}n_{k},$ (9)
represents solute interactions with grain boundaries and corresponds to the term $\sum_{k}J_{sg}\xi_{k}\phi(n_{k})$ in Eq.(4). Direct mapping in not possible because Eq.(9) represents interactions of solute atoms with all sites that have wrong neighbors ($n_{k}\neq 0$), whereas the term $\sum_{k}J_{sg}\xi_{k}\phi(n_{k})$ includes the function $\phi(n_{k})$ specifically targeting grain boundary environments. We can only say that $J_{sg}$ is physically related to the energy difference $(\varepsilon_{AA}^{\prime}-\varepsilon_{AB}^{\prime})-(\varepsilon_{AA}-\varepsilon_{AB})$.
The term
$\dfrac{1}{2}\sum_{k}\sum_{(k,l)}\xi_{k}\xi_{l}(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB}),$
describes solute-solute interactions and is similar to the term $\tfrac{1}{2}\sum_{k}\sum_{(kl)}\xi_{k}\xi_{l}J_{ss}$ in Eq.(4). This leads to the mapping
$J_{ss}=(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB}).$ (10)
The last term in Eq.(5),
$\dfrac{1}{2}\sum_{k}\sum_{(k,l)}m_{kl}\xi_{k}\xi_{l}\left[(\varepsilon_{BB}^{\prime}+\varepsilon_{AA}^{\prime}-2\varepsilon_{AB}^{\prime})-(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB})\right],$ (11)
is similar to the term $\tfrac{1}{2}\sum_{k}\sum_{(kl)}\xi_{k}\xi_{l}J_{ssg}\phi(n_{k,l})$ in Eq.(4). However, direct mapping is again impossible. Eq.(11) accounts for the difference in solute-solute interactions inside the perfect lattice and between solute atoms occupying sites with wrong neighbors, whereas the Potts model makes this effect specific to grain boundaries. Furthermore, the Potts model considers the average “wrongness” $n_{k,l}$ of the two sites. All we can say is that the parameter $J_{ssg}$ is physically related to the change in the interaction parameter $(\varepsilon_{BB}+\varepsilon_{AA}-2\varepsilon_{AB})$ when the solute is at grain boundaries.
It is instructive to reformulate the above equations in terms of the regular solution parameter
$\Omega=\varepsilon_{AB}-\dfrac{\varepsilon_{BB}+\varepsilon_{AA}}{2}.$ (12)
The solute-solute interaction factor (10) becomes
$J_{ss}=-2\Omega,$ (13)
while Eq.(11) takes the form
$\sum_{k}\sum_{(k,l)}m_{kl}\xi_{k}\xi_{l}(\Omega^{\prime}-\Omega).$ (14)
Figure 4: Schematic grain boundary separating perfectly crystalline grains. The solid red segments represent boundaries between sites with different orientations. The two black circles symbolize solute atoms segregated at the grain boundary.
6 Application to a single planar grain boundary
In this section, we apply the general mapping relations established in the previous section to a particular case of a single planar grain boundary shown schematically in Figure 4. The grains are considered perfectly single-crystalline ($n_{k}=0$). In addition, we will specialize the above equations in the constraint $J_{s}=0$, which we adopted in this work to simplify the parametric study. Eq.(8) gives $\varepsilon_{AB}-\varepsilon_{AA}=0$. Considering that we have shifted the reference energy to achieve $\varepsilon_{AA}=0$, we arrive at $\varepsilon_{AB}=\varepsilon_{AA}=0$. Thus, there is only one independent bond energy: $\varepsilon_{BB}$.
For the grain boundary geometry shown in Figure 4, all boundary sites have $n_{k}=2$ and thus $\phi(n_{k})=1$. Each boundary site has excess energy $J_{gg}$ and occupies the grain boundary area $A=\sqrt{2}a^{2}$, where $a$ is the side of the square that represents the site. As a result, the grain boundary energy per unit area in a solute-free system is
$\gamma=\dfrac{J_{gg}}{A}=\dfrac{J_{gg}}{\sqrt{2}a^{2}}.$ (15)
Eq.(9) predicts that the energy change in moving a single solute atom from inside a grain to the grain boundary is
$E_{s}=2(\varepsilon_{AB}^{\prime}-\varepsilon_{AA}^{\prime})-2(\varepsilon_{AB}-\varepsilon_{AA})=2(\varepsilon_{AB}^{\prime}-\varepsilon_{AA}^{\prime}),$ (16)
whereas the respective term $\sum_{k}J_{sg}\xi_{k}\phi(n_{k})$ in Eq.(4) reduces to $J_{sg}$. This allows us to identify $J_{sg}$ as the grain boundary segregation energy:
$E_{s}=J_{sg}.$ (17)
According to Eq.(13),
$\Omega=-\dfrac{J_{ss}}{2}.$ (18)
To identify $J_{ssg}$, we note that the term $\dfrac{1}{2}\sum_{k}\sum_{(kl)}\xi_{k}\xi_{l}J_{ssg}\phi(n_{k,l})$ in Eq.(4) simplifies to $\tfrac{1}{2}\sum_{k}\sum_{(kl)}\xi_{k}\xi_{l}J_{ssg}$, while Eq.(14) becomes
$\sum_{k}\sum_{(k,l)}\xi_{k}\xi_{l}(\Omega^{\prime}-\Omega).$ (19)
It follows that
$(\Omega^{\prime}-\Omega)=-\dfrac{1}{2}J_{ssg}.$ (20)
Equations (15), (17), (18), and (20) express the experimentally relevant properties $\gamma$, $E_{s}$, $\Omega$, and $\Omega^{\prime}$ through model parameters $J_{gg},$ $J_{sg}$, $J_{ss}$, and $J_{ssg}$, respectively.
7 Numerical estimates
In the KMC simulations reported in the paper, we used the dimensionless variables by normalizing the temperature and $J$-factors by $J_{gg}$, see Table 1 in the main text. Inverting Eq.(15), we obtain
$J_{gg}=\gamma\sqrt{2}a^{2},$ (21)
indicating that $J_{gg}$ can be interpreted as the energy per atom in the grain boundary core.
For numerical estimates, we will use the typical grain boundary energy $\gamma=1\>\mathrm{J}/\mathrm{m}^{2}$ and estimate $a$ by 1 nm. Eq.(21) then gives us $J_{gg}=0.88$ eV/atom, which is a physically reasonable number for the excess energy of a grain boundary atom. Using this value of $J_{gg}$, we calculated the energies $J_{sg}$, $J_{ss}$, and $J_{ssg}$ and then obtained the thermodynamic properties $E_{s}$, $\Omega$, and $\Omega^{\prime}$ for the three cases for which the phase diagrams were computed in the paper.
Case 1 Case 2 Case 3
$J_{sg}/J_{gg}$ 1.6-1.6 1.6-1.6 1.6-1.6
$J_{ss}/J_{gg}$ 0 0.25-0.25 $0.25$
$J_{ssg}/J_{gg}$ 0 0.05-0.05 $0.05$
$E_{s}$ (eV/atom) 1.41-1.41 1.41-1.41 1.41-1.41
$E_{s}$ (kJ/mol) 136-136 136-136 136-136
$\Omega$ (eV/atom) 0 $0.110$ 0.110-0.110
$\Omega$ (kJ/mol) 0 $10.61$ 10.61-10.61
$\Omega^{\prime}$ (eV/atom) 0 $0.132$ 0.132-0.132
$\Omega^{\prime}$ (kJ/mol) 0 $12.73$ 12.73-12.73
$H_{\mathrm{mix}}$ (kJ/mol) 0 $50.94$ 50.94-50.94
Table 3: Potts model parameters normalized to $J_{gg}$ are compared with thermodynamic properties: grain boundary segregation energy $E_{s}$, the regular solution parameters inside the grains ($\Omega$) and in grain boundaries ($\Omega^{\prime}$), and the dilute heat of mixing ($H_{\mathrm{mix}}$). The cases 1, 2, and 3 correspond to parameter sets used for computing the phase diagrams shown in Figures 3, 6 and 7 of the main text.
Table 3 summarizes the results in both eV/atom and practical units kJ/mol. In addition to the regular solution parameter $\Omega$, the table includes the perfect-lattice heat of mixing $H_{\mathrm{mix}}=z\Omega$ in the dilute limit.
The predicted magnitudes of $E_{s}$, $\Omega$, $\Omega^{\prime}$, and $H_{\mathrm{mix}}$ are well within the range of typical values reported in previous publications [11, 5, 69, 6, 68, 9]. The notable feature of the present simulations is that the stable polycrystalline structure was found in alloys with a negative heat of mixing, corresponding to the ordering trend (Case 3 in Table 3). Most of the previous calculations [11, 5, 69, 6, 68, 9] considered phase-separating alloys with $H_{\mathrm{mix}}>0$. However, recent publications [68, 7] extended the previous analyses to alloys with $H_{\mathrm{mix}}<0$, and stable nanocrystalline states were also observed. For example, Figure 6 in [68] presents a diagram showing a nanocrystalline stability domain at $\Omega^{\prime}<\Omega<0$. Furthermore, the values of $\Omega$ and $\Omega^{\prime}$ of the representative point g in this domain are comparable to the values for Case 3 in Table 3. Finally, the temperature interval of the stable polycrystalline state in Case 3 is approximately between 1000 and 3000 K. Although these temperatures are higher than one could expect, they are not unreasonable given the assumptions and approximations underlying this calculation.
A more detailed comparison with previous publications [11, 5, 69, 6, 68, 9] is complicated due to intrinsic differences between models. In addition, in both models, different constraints are imposed to reduce the dimensionality of the parameter space. For example, we imposed the constraint $\varepsilon_{AB}=\varepsilon_{AA}=0$, whereas the previous models [11, 5, 69, 6, 68, 9] imposes the constraints (in our notation) $\varepsilon_{AA}=\varepsilon_{BB}$ and $\varepsilon_{AA}^{\prime}=\varepsilon_{BB}^{\prime}$.